Louisiana State University Health Sciences Center Administration & Finance
 

Friday, August 18, 2006

  • Reminder: The contractor working on 1542 Tulane Avenue will need to set up cranes in the Lakeside Lot beginning early Monday morning, August 21st. The lot will be closed Monday and Tuesday, August 21st and 22nd. No vehicles will be allowed in the lot during this process and all vehicles must be removed prior to 5:00 AM Monday, August 21st.
  • Of the windows that were originally identified as having been damaged by the hurricane on the downtown campus, 96 window systems have been replaced.
    • Three window systems are still on order.
    • There are three new windows that are included in a proposed change order that will require approximately 8 weeks to be delivered, once the change is approved.
    • The six damaged windows at 1542 Tulane, which were part of the original project, remain to be replaced.
    • An additional change order is being prepared for submission to the architects. This will address leaks in the atrium glass ceiling in the Resource Center. The revised proposal including all new items will be sent to the architect Monday.
  • The repairs to building systems at 1542 Tulane are progressing.
    • Electrical panels PL2 and 3 were mounted today.
    • The PL1 cables have been run but are not yet terminated. The contractor is in the process of building the cable supports from the ceiling of the basement.
    • Several men are working on tying in the feeder cables from the 1KV transformer to the 3KV switchgear which powers the elevators.
    • The 600 Amp Disconnect is mounted and the feeder cables are terminated to it.
    • Over the coming weekend the troughs will be mounted for the cables that power the sump pumps.
  • Work in the basement of the Dental School continued today.
    • The insulation of the water piping in the Administration building continued.
    • Mold remediation has begun in some of the upper rooms that were heavily impacted; Rooms 6413B, 6409, & 7409A.
    • The piping for repairing the medical gas system has been delivered and is staged onsite.
    • Replacement of governor cables for Elevator #4 in the Clinic Bldg underway. Elevator doors for Elevator #4 on the 1st floor has been removed...awaiting new replacement doors.
    • The removal of the flood damaged medical records is approximately 50% complete.
    • The new natural gas piping has been completed, but it has been shut off until the system testing can be scheduled and completed.
    • The recently emergency chiller failed today during a brief afternoon thunderstorm. The 24 hour response team was called to bring the chiller back online.
  • In the MEB cafeteria, the contractor is installing the grill hood exhaust fan which required repairs following the storm.